Solve the system of equations using substitution
3x + 4y=12
2x + y=10

Found 2 solutions by angelicnwa, mananth:
Answer by angelicnwa(2)

again, I am not a math person, but I would think you would plug in the numbers.
3(4) + 4(0) = 12 12+0=12
2(4) + 2 = 10 8+2=10
Just a guess

Answer by mananth(16946)

2 x + 1 y = 10 .............1
1 y = 10 + -2 x
/ 1 =
y = 10 + -2 x
3 x + 4 y = 12 .....................2
Substitute y in (2)
3x+4(10-2x)= 12
3x+40-8x= 12
3x-8x =-28
-5x = -28
/ -5
x= 28/ 5
Plug the value of x in (1)
2 x + 1 y = 10
2 * 28/ 5 + 1 y = 10
11 1/ 5 + 1 y = 10
1 y = -6/ 5
/ 1
y= -6/ 5
(28/5 , -6/5)
m.ananth@hotmail.ca